OpenStudy (anonymous):

solve 5x^2-x+4=0 by completing the square.

myininaya (myininaya):

5x^2-x=-4 x^2-1/5*x=-4/5 x^2-1/5*x+(1/[2*5])^2=-4/5+(1/[2*5])^2 x^2-1/5*x+(1/10)^2=-4/5+(1/10)^2 (x-1/10)^2=-4/5+1/100 do u think u can finish it

myininaya (myininaya):

(x-1/10)^2=-4/5*20/20+1/100 (x-1/10)^2=-80/100+1/100 (x-1/10)^2=-79/100

myininaya (myininaya):

mome do you have any question you just take the square root of both sides and then add 1/10 and you are done you might be able to simplify and stuff you can say sqrt{-1} is i

OpenStudy (anonymous):

myininaya, I was confused because I thought you take half of the middle term (x) or 1.

OpenStudy (anonymous):

\[\frac{5x^2-x+4}{5}=\frac{0}{5}\]

myininaya (myininaya):

you do mome 1/2*1/5 and then don't forget to square after you take a half of the middle term

OpenStudy (anonymous):

\[x^2-\frac{1}{5}x+\frac{4}{5}=0\]

OpenStudy (anonymous):

\[\frac{\frac{1}{5}}{2}=\frac{1}{10}\]

OpenStudy (anonymous):

\[(\frac{1}{10})^2=\frac{1}{100}\]

OpenStudy (anonymous):

\[(x-\frac{1}{10})^2=-\frac{4}{5}+\frac{1}{100}\]can u solve from now?

OpenStudy (anonymous):

still trying to understand it. California

OpenStudy (anonymous):

let me know when u ready go next step

OpenStudy (anonymous):

ready

OpenStudy (anonymous):

frist step divi 5 get rid coefficient second step take (1/5)/2 = 1/10 (x- 1/10 )^2 thrid step (1/10)^=1/100 move 4/5 to right hand side u get (-4/5 )+ 1/100

OpenStudy (anonymous):

now get rid with square on the left , u need square root on the right hand side

OpenStudy (anonymous):

\[x-\frac{1}{10}=\pm \sqrt{\frac{-80+1}{100}}\]

OpenStudy (anonymous):

\[x-\frac{1}{10}=\pm \sqrt{\frac{-79}{100}}\]

OpenStudy (anonymous):

\[x=\frac{1}{10}\pm\frac{\sqrt{-1}*\sqrt{79}}{10}\]

OpenStudy (anonymous):

\[x=\frac{1}{10}\pm i \frac{\sqrt{79}}{10}\]

OpenStudy (anonymous):

is your answer

OpenStudy (anonymous):

?